Mar 31, 2021 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

  1. Short term CALWG Items
  2. Long term items
  3. Calorimeter work packages

Tentative Agenda

  1. Announcements
  2. Run Preparations
    1. FCAL
    2. BCAL
  3. Calibration
    1. BCAL
      1. Summer 2020 Calibration NonlinearCorrections (ksuresh)
  4. Monitoring
    1. BCAL (offline)
      1. Tolga - Big update batches 10-12 [2]
    2. FCAL (offline)
      1. RunPeriod-2019-11
        1. Batch 10-12 monitoring launch ver26
        2. Batch 1 LED efficiencies and LED time shifts
        3. comparison to raw data in progress
  5. Reconstruction
  6. Simulations
  7. Action Items
  8. Any other business


Attending: Elton, Jon, Varun, Sean, Tolga, Colin, Susan, Karthik, Churamani, Matt, Mark, Sasha, Malte, Zisis

  1. Announcements: none.
  2. Run Preparations
    1. FCAL: Matt reported that Paul Smith highlighted that failures of repaired bases sometimes are a result of collateral damage when removing the conformal coating. The repairs will continue under the current method, but there is thought of melting the coating in future tests. This info has been passed on to Mark, Chris and Nick. Colin recommended that the FCAL should be powered up for at least 2 months before the run; everyone agreed. Colin also suggested that the bases be tested with LEDs as sometimes the board connection is not good (related to improper routing of the signal cable); Jon added that this can lead to charge accumulation and discharge that affects nearby bases. The bases have been inserted in the FCAL though, so it will be powered up in its dark room and tested.
    2. BCAL: no report.
  3. Calibration
    1. BCAL: A summer 2020 calibration update was given by Karthik on the non-linear energy correction. Events over 3.5 GeV are ignored (stats are too low). Change: at 11 GeV the curve is now anchored to 1.0 (previously it was 0.95), to avoid overcorrecting. He will investigate the pixel/count factor in the coming weeks. Batches 9-12 were treated as a single batch, due to their similarities. Runs are still missing; resubmission is only 10% successful. Sean asked that Karthik resubmits and will work with him offline to diagnose the issue. The calibration looks very good: the pi0 mass values are within 0.2% of the expected mass; constants are still in Karthik's local ccdb, but they have not changed much since the previous set of constants.
  4. Monitoring
    1. BCAL (offline)
      1. Tolga gave an update on batches 10-12. Things quite good in batch 10, worse in batch 12. A few hot spots, near modules 34/35. These hot spikes sometimes disappear and then reappear. Mark will think about a more robust pedestal calibration as the FADCs drift with temperature. Discussion continued about module 37. Varun reminded us of the issue reported in the fall: E-log.
    2. FCAL (offline): Susan continued her work on the FCAL monitoring, for batches 10-12. Mostly routine. Was slowed down by tape issues (sub files). Showed a plot clearly demonstrating the 2ns timing shift that Sean had talked about. Sean will try to correlate these shifts with any changes in the calibration of those runs.
  5. Reconstruction. Churamani reported on his photon efficiency studies using the η' tag and probe method. He showed the efficiency which looks like a nice result. Mark suggested that he adds a plot of efficiency versus energy to complement the plot versus θγ. The MC seems to stop after 31 degrees. Sean asked for a table of the results and cautioned about using data below the PS range: care must be taken to match MC and data. Some discussion ensued on the statistics of this method. With the looser energy cut, there are about 2400 events now.
  6. Simulations: Tolga plans to return to his neutron ID work in the next couple of weeks.
  7. Action Items: Zisis added material to the BCAL Wiki and updated the short-term tasks.
  8. Any other business: none.